A ball falls 20meters from rest to the ground below

a)What is the velocity as it strikes the ground?

b)How long does it take to fall?


1
Expert's answer
2022-06-03T15:25:40-0400

(a)

"v=\sqrt{2gh}=\sqrt{2*9.8*20}=19.8\:\rm m/s"

(b)

"t=\frac{v}{g}=\frac{19.8}{9.8}=2.0\:\rm s"